Corbel Corporation has two divisions: Division A and Division B. Last month, the company reported a contribution margin of $44,300 for Division A. Division B had a contribution margin ratio of 40% and its sales were $232,000. Net operating income for the company was $32,600 and traceable fixed expenses were $55,800. Corbel Corporation's common fixed expenses were:

  • $48,700

  • $55,800

  • $104,500

  • $137,100
Answer #1

Total Contribution margin

= 44,300+(40%*232,000) = 137,100

Segment margin = Total Contribution margin - Traceable fixed expenses

= 137,100 - 55,800 = 81,300

Net operating income = Segment margin - Common fixed expenses

32,600 = 81,300 - Common fixed expenses

Common fixed expenses = 81,300 - 32,600

= 48,700